# Quantum mixing on large Schreier graphs

### v1

@@ Appendix A Strong convergence in distribution
It coincides with the operator norm of $\lambda(a)$ in $\ell^{2}(\Gamma)$.
@@ Appendix B Rapid decay property
Assume that $\Gamma$ has property RD with constant $C_{1}>0$ as in Definition 1.4 and let $S$ be a finite set of generators of $\Gamma$. Then for all $C^{\prime}_{1}>2C_{1}+1$, there exist a constant $C>0$ such that for all $p\in\mathbb{C}[\Gamma]$,
@@
For integer $n\geq 0$, let $p_{n}=\sum_{g:|g|=n}p_{g}g$. The property RD implies that, for some $C>0$,
@@
The next lemma extracted from [58] (see also [74, Lemma 2.26]) shows that to prove strong convergence is enough to consider Markovian $p\in\mathbb{C}[\Gamma]$, that is $p\in\mathbb{R}[\Gamma]$ such that $\sum_{g}p_{g}=1$ and $p_{g}\geq 0$ for all $g\in\Gamma$. We start by introducing a weakening of the rapid decay property, satisfied by all groups with subexponential growth:
@@
We say that $\Gamma$ has the weak rapid decay (RD’) property if for some subexponential function $E$, for any $p\in\mathrm{C}[\Gamma]$,
@@
We may write $p\in\mathbb{C}[\Gamma]$ as $p=p_{1}-p_{2}+ip_{3}-ip_{4}$ with for $i\in\{1,\ldots,4\}$, $p_{i}\in\mathbb{R}(\Gamma)$ and for all $g\in\Gamma$ $(p_{i})_{g}\geq 0$,
@@
Since $\|p\|_{2}=\|\lambda(p)\delta_{e}\|_{2}\leq\|\lambda(p)\|_{\rm op}$ the conclusion follows from Lemma A.2. ∎
The next proposition asserts that for amenable groups of subexponential growth convergence in distribution implies strong convergence. We refer also to Magee [57, Section 2.1] for a closely related discussion.
@@
Let $\Gamma$ be a group of subexponential growth. Let $\rho_{N}$ be a sequence of unitary representation of $\Gamma$ of dimension $N$ (along a subsequence of integers). If $(\rho_{N})$ converges in distribution toward $\lambda$ and $H_{N}$ is an invariant subspace of dimension $o(N)$ then $({\rho_{N}}_{|H_{N}^{\perp}})$ converges strongly in distribution toward $\lambda$.
@@
Due to Proposition A.1, we only need to show the inequality (A.3) for any $p\in\mathbb{C}[\Gamma]$. To prove it, we may further assume by Lemma B.3 (recalling that $\Gamma$ has the RD’ property since it has subexponential growth) that $p\in\mathbb{C}[\Gamma]$ is Markovian. Let us check that $\|\lambda(p)\|_{\rm op}=\sum_{g}p_{g}=1$. We have $\|\lambda(p)\|_{\rm op}\leq\|p\|_{1}=1$. Conversely, since all groups of subexponential growth are amenable, there exists a Følner sequence $(F_{k})$ for $\Gamma$. If $\chi_{k}$ denotes the $\ell^{2}$-normalized characteristic function of $F_{k}$, then, for any $g\in\Gamma$,
@@ Appendix C Convergence rate
Let $\mu,\nu$ be two probability measures on $a,a]\subset\mathbb{R}$ and $n\in\mathbb{N}$. Assume that $\mu$ has a density bounded by $b$ on an interval $I\subset[-a,a]$ and that for all integers $k\in\left[2n\right]$,
@@
[…]
@@
where $p$ is any polynomial of degree at most 2(n1)2(n-1) such that $p(\lambda)\geq 0$ on [a,a][-a,a] and $p(t)\geq 1$ (the condition $p(\lambda)\geq 0$ on $\mathbb{R}$ is required in [42] but non-negativity on [a,a][-a,a] is sufficient since all roots of orthogonal polynomials are in the support of their associated measure).
@@
[…]
@@
where $C=\|p\|_{1}\left(3{C_{0}}+\|p\|_{1}/(n\epsilon^{2})\right)$ and $\Lambda_{J}$ is the set of eigenvalues of $\rho_{N}(p)$ in $J$ (counting multiplicities).
@@
Set $P_{N}=\rho_{N}(p)\in M_{N}(\mathbb{C})$ and let $\mu_{P_{N}}$ be the empirical distribution of eigenvalues of $P_{N}$. We have $|\Lambda_{J}|=N\mu_{P_{N}}(J)$. Also, for $\psi\in\mathbb{C}^{N}$, recall the definition of $\mu_{P_{N}}^{\psi}$ in (3.6). As explained in the proof of Lemma 3.2, if $x\notin{\rm Bad}_{S}(n)$ then for all integers $k\in\left[2n\right]$,
$\int\lambda^{k}d\mu^{\delta_{x}}_{P_{N}}=\int\lambda^{k}d\mu^{\delta_{x}}_{p}.$
The support of $\mu_{p}$ and $\mu^{\delta_{x}}_{P_{N}}$ is contained in [a,a][-a,a] with $a=\|p\|_{1}=\sum_{g}|p_{g}|$. The conclusion follows from Proposition C.1, together with (3.7). ∎
We may now discuss the rates of convergence that we can obtain from the proofs of Theorem 1.5 and Theorem 1.7. A similar discussion can be done for Theorem 1.3. For ease of discussion, we will restrict ourselves to the following quantified convergence in distribution of $\rho_{N}$ toward $\lambda$: for some $\alpha>0$, $\beta\in[0,1)$, $C>0$
@@
where we recall that ${\rm Bad}_{S}(r)=\mathrm{Fix}^{*}_{N}(B_{S}(2r))$. This situation occurs for example for congruence actions of linear groups (for example $\Gamma=\mathrm{SL}_{d}(\mathbb{Z})$ and $\rho_{N_{n}}(g)=g\mod[n]$ acting on $\mathrm{SL}_{d}(\mathbb{Z}/n\mathbb{Z})$) then (C.2) holds for some $\alpha>0$ and $|{\rm Bad}_{S}(r_{N})|=0$ (for $d=2$ and $n$ prime, see [41, Eqn (2.6)], the argument holds for any $d$ and $n$ non-prime). Also, this situation is typical in random actions of free groups (see for example [22, Appendix D]), the argument could also be extended to random actions free products of finite groups or surface groups, see [66] and references therein.
We will also assume for some $C>0$, $C^{\prime}_{1}>2C_{1}+1$ where $C_{1}$ is as in (1.13) and $\gamma\in(0,1]$, $z=\lambda+i\eta$, $\lambda\in I$,
@@
Assume that $\Gamma$ has property RD with constant $C_{1}$ and that $\rho_{N}$ satisfies (C.2)-(C.4). Let $p\in\mathbb{C}[\Gamma]$ such that $p=p^{*}$ and an interval $I_{0}$ such that (1.5)-(C.3) holds for $P=\lambda(p)$. Let $I$ be a closed interval in the interior of $I_{0}$ and $(\varphi_{\alpha})$ be an orthonormal eigenbasis $P_{N}=\rho_{N}(p)$. Then, for any finite $T\subset\Gamma$, there exists $C>0$ such that for any $T$-local matrix $(K_{N})$ in $M_{N}(\mathbb{C})$ with $\|K_{N}\|_{1,\infty}\leq 1$, we have for all $E_{1},E_{2}\subset I$,
@@
Also, by Corollary C.2 and (1.5), assumption (C.2) implies that for some $C>0$, $1/C\leq|\Lambda_{J}|/(N|J|)\leq C$ for all intervals of $J\subset I$ of length $|J|\geq C/\ln N$. If (C.4) holds, in the proof of Theorem 1.7 we take
$\eta_{N}=\frac{a\ln\ln N}{\ln N},$
$\epsilon_{N}=1/(2\ln(\max(1/\eta_{N},4)))\sim 1/(2\ln\ln N)$ and $a>(2c)/\alpha$ chosen so that $n_{N}=\lceil c/(\eta_{N}\epsilon_{N})\rceil\sim(2c/a)\ln N\leq r_{N}$ for all $N$ large enough. Then using $|{\rm Bad}_{S}(n_{N})|\leq CN^{\beta}$ from (C.2), we readily deduce the requested bound from (4.8) and Lemmas 2.1-2.2.
@@
$\epsilon_{N}=1/(2\ln(\max(1/\eta_{N},4)))\sim 1/(2\ln\ln N)$ and $n_{N}=\lceil c/(\eta_{N}\epsilon_{N})\rceil\sim(2c/a)\ln N/\ln\ln N$ with $c>0$ as in Lemma 2.3 and $\ell_{N}=\lfloor r_{N}/n_{N}\rfloor\sim(a\alpha/(2c))\ln\ln N$. For $a$ large enough, the conditions (4.14) are satisfied. Further, for any $b>0$, we may find $a>0$ large enough, so that the dimension of the vector space $H_{N}$ defined in the proof of Theorem 1.5 has dimension at most $N/(\ln N)^{b}$. The conclusion follows similarly. Note that it is also possible to pick a larger $\eta_{N}\to 0$ to obtain a vector space $H_{N}$ of smaller dimension. ∎

### v2

@@ Appendix A Strong convergence in distribution
It coincides with the operator norm of $\lambda(a)$ in $\ell^{2}(\Gamma)$. Here $\tau(a)=a_{e}$.
@@ Appendix B Rapid decay property
Assume that $\Gamma$ has the RD property with constant $C_{1}>0$ as in Definition 1.4 and let $S$ be a finite set of generators of $\Gamma$. Then for all $C^{\prime}_{1}>2C_{1}+1$, there exist a constant $C>0$ such that for all $p\in\mathbb{C}[\Gamma]$,
@@
For integer $n\geq 0$, let $p_{n}=\sum_{g:|g|=n}p_{g}g$. The RD property implies that, for some $C>0$,
@@
The next lemma extracted from [58] (see also [74, Lemma 2.26]) shows that to prove strong convergence, it is enough to consider Markovian $p\in\mathbb{C}[\Gamma]$, that is $p\in\mathbb{R}[\Gamma]$ such that $\sum_{g}p_{g}=1$ and $p_{g}\geq 0$ for all $g\in\Gamma$. We start by introducing a weakening of the rapid decay property, satisfied by all groups with subexponential growth:
@@
We say that $\Gamma$ has the weak rapid decay (RD’) property if for some subexponential function $E$, for any $p\in\mathbb{C}[\Gamma]$,
@@
We may write $p\in\mathbb{C}[\Gamma]$ as $p=p_{1}-p_{2}+ip_{3}-ip_{4}$ with for $i\in\{1,\ldots,4\}$, $p_{i}\in\mathbb{R}[\Gamma]$ and for all $g\in\Gamma$ $(p_{i})_{g}\geq 0$,
@@
Since $\|p\|_{2}=\|\lambda(p)\delta_{e}\|\leq\|\lambda(p)\|_{\rm op}$ the conclusion follows from Lemma A.2. ∎
The next proposition asserts that for amenable groups of subexponential growth, convergence in distribution implies strong convergence. We refer also to Magee [57, Section 2.1] for a closely related discussion.
@@
Let $\Gamma$ be a group of subexponential growth. Let $\rho_{N}$ be a sequence of unitary representations of $\Gamma$ of dimension $N$ (along a subsequence of integers). If $(\rho_{N})$ converges in distribution toward $\lambda$ and $H_{N}$ is an invariant subspace of dimension $o(N)$, then $({\rho_{N}}_{|H_{N}^{\perp}})$ converges strongly in distribution toward $\lambda$.
@@
Due to Proposition A.1, we only need to show the inequality (A.3) for any $p\in\mathbb{C}[\Gamma]$. To prove it, we may further assume by Lemma B.3 (recalling that $\Gamma$ has the RD’ property since it has subexponential growth) that $p\in\mathbb{C}[\Gamma]$ is Markovian. Let us check that $\|\lambda(p)\|_{\rm op}=\sum_{g}p_{g}=1$. We have $\|\lambda(p)\|_{\rm op}\leq\|p\|_{1}=1$. Conversely, since all groups of subexponential growth are amenable, there exists a Følner sequence $(F_{k})$ for $\Gamma$. If $\chi_{k}$ denotes the $\ell^{2}$-normalized characteristic function of $F_{k}$, then for any $g\in\Gamma$,
@@ Appendix C Convergence rate
Let $\mu,\nu$ be two probability measures on $a,a]\subset\mathbb{R}$ and $n\in\mathbb{N}$. Assume that $\mu$ has a density bounded by $b$ on an interval $I\subset[-a,a]$ and that for all integers $k\leq 2n$,
@@
Then, for any $t\in I$ at distance at least $\epsilon>0$ from $\mathbb{R}\setminus I$, we have
@@
where $p$ is any polynomial of degree at most 2(n1)2(n-1) such that $p(\lambda)\geq 0$ on [a,a][-a,a] and $p(t)\geq 1$ (the condition $p(\lambda)\geq 0$ on $\mathbb{R}$ is required in [42], but non-negativity on [a,a][-a,a] is sufficient since all roots of orthogonal polynomials are in the support of their associated measure).
@@
Let $p\in\mathbb{C}[\Gamma]$ with $p=p^{*}$ and let $S\subset\Gamma$ be its support. For integer $n\geq 1$, let ${\rm Bad}_{S}(n)=\mathrm{Fix}_{N}^{*}(B_{S}(2n))$ where $B_{S}(n)$ is the ball of radius $n$ in $\mathrm{Cay}(\Gamma,S)$. If $I$ is an interval and $\mu_{p}$ admits a density on $I$ uniformly bounded by $C_{0}$}, then for any interval $J\subset I$ at distance at least $\epsilon>0$ from $\mathbb{R}\setminus I$ we have
@@
where $C=2\|p\|_{1}\left(3{C_{0}}+\|p\|_{1}/(n\epsilon^{2})\right)$ and $\Lambda_{J}$ is the set of eigenvalues of $\rho_{N}(p)$ in $J$ (counting multiplicities).
@@
Set $P_{N}=\rho_{N}(p)\in M_{N}(\mathbb{C})$ and let $\mu_{P_{N}}$ be the empirical distribution of eigenvalues of $P_{N}$. We have $|\Lambda_{J}|=N\mu_{P_{N}}(J)$. Also, for $\psi\in\mathbb{C}^{N}$, recall the definition of $\mu_{P_{N}}^{\psi}$ in (3.6). As explained in the proof of Lemma 3.2, if $x\notin{\rm Bad}_{S}(n)$ then for all integers $k\leq 2n$,
$\int\lambda^{k}d\mu^{\delta_{x}}_{P_{N}}=\int\lambda^{k}d\mu_{p}.$
The support of $\mu_{p}$ and $\mu^{\delta_{x}}_{P_{N}}$ is contained in [a,a][-a,a] with $a=\|p\|_{1}=\sum_{g}|p_{g}|$. Hence
$|\Lambda_{J}|=N\mu_{P_{N}}(J)\geq\sum_{x\notin{\rm Bad}_{S}(n)}\mu_{P_{N}}^{\delta_{x}}(J)\geq(N-|{\rm Bad}_{S}(n)|)\left(\mu_{p}(J)-\frac{C}{n}\right)$
where we used (3.7) and then Proposition C.1. For the right-hand side we use $\mu_{P_{N}}^{\delta_{x}}(J)\leq 1$ and Proposition C.1 to get
$|\Lambda_{J}|=N\mu_{P_{N}}(J)=\sum_{x\in{\rm Bad}_{S}(n)}\mu_{P_{N}}^{\delta_{x}}(J)+\sum_{x\notin{\rm Bad}_{S}(n)}\mu_{P_{N}}^{\delta_{x}}(J)\leq N\left(\mu_{p}(J)+\frac{C}{n}\right)+|{\rm Bad}_{S}(n)|$
which concludes the proof. ∎
We may now discuss the rates of convergence that we can obtain from the proofs of Theorem 1.5 and Theorem 1.7. Similar arguments can be given for Theorem 1.3. For ease of discussion, we will restrict ourselves to the following quantified convergence in distribution of $\rho_{N}$ toward $\lambda$: for some $\alpha>0$, $\beta\in[0,1)$, $C>0$
@@
where we recall that ${\rm Bad}_{S}(r)=\mathrm{Fix}^{*}_{N}(B_{S}(2r))$. This situation occurs for example for congruence actions of linear groups (for example $\Gamma=\mathrm{SL}_{d}(\mathbb{Z})$ and $\rho_{N_{n}}(g)=g\mod[n]$ acting on $\mathrm{SL}_{d}(\mathbb{Z}/n\mathbb{Z})$) then (C.2) holds for some $\alpha>0$ and $|{\rm Bad}_{S}(r_{N})|=0$ (for $d=2$ and $n$ prime, see [41, Eqn (2.6)], the argument holds for any $d$ and $n$ non-prime). Also, this situation is typical in random actions of free groups (see e.g. [22, Appendix D]), the argument could also be extended to random actions of free products of finite groups or surface groups, see [66] and references therein.
We will also assume for some $C>0$, $C^{\prime}_{1}>2C_{1}+1$, where $C_{1}$ is as in (1.13), and $\gamma\in(0,1]$, that if $z=\lambda+i\eta$, $\lambda\in I$, then
@@
Assume that $\Gamma$ has the RD property with constant $C_{1}$ and that $\rho_{N}$ satisfies (C.2)-(C.4). Let $p\in\mathbb{C}[\Gamma]$ such that $p=p^{*}$ and an interval $I_{0}$ such that (1.5)-(C.3) holds for $P=\lambda(p)$. Let $I$ be a closed interval in the interior of $I_{0}$ and $(\varphi_{\alpha})$ be an orthonormal eigenbasis for $P_{N}=\rho_{N}(p)$. Then, for any finite $T\subset\Gamma$, there exists $C>0$ such that for any $T$-local matrix $(K_{N})$ in $M_{N}(\mathbb{C})$ with $\|K_{N}\|_{1,\infty}\leq 1$, we have for all $E_{1},E_{2}\subset I$,
@@
Also, by Corollary C.2 with $n:=r_{N}$ and (1.5), assumption (C.2) implies that for some $C>0$, $1/C\leq|\Lambda_{J}|/(N|J|)\leq C$ for all intervals of $J\subset I$ of length $|J|\geq C/\ln N$. If (C.4) holds, in the proof of Theorem 1.7 we take $\eta$ depending on $N$, namely
$\eta_{N}=\frac{a\ln\ln N}{\ln N}$
where $a>(2c)/\alpha$ with $c$ as in Lemma 2.3 and $\alpha$ as in (C.2). Let $\epsilon_{N}=1/(2\ln(\max(1/\eta_{N},4)))\sim 1/(2\ln\ln N)$, and observe that $n_{N}=\lceil c/(\eta_{N}\epsilon_{N})\rceil\sim(2c/a)\ln N\leq r_{N}-1$ for all $N$ large enough. Then using $|{\rm Bad}_{S}(n_{N}+1)|\leq CN^{\beta}$ from (C.2), we readily deduce the requested bound from (4.8) and Lemmas 2.1-2.2.
@@
$\epsilon_{N}=1/(2\ln(\max(1/\eta_{N},4)))\sim 1/(2\ln\ln N)$ and $n_{N}=\lceil c/(\eta_{N}\epsilon_{N})\rceil\sim(2c/a)\ln N/\ln\ln N$ with $c>0$ as in Lemma 2.3 and $\ell_{N}=\lfloor r_{N}/n_{N}\rfloor\sim(a\alpha/(2c))\ln\ln N$. Further, in the proof of Theorem 1.5, instead of (4.17), given any sequence $(g_{N})$ with $g_{N}\rightarrow+\infty$ as $N\rightarrow+\infty$, we may impose that the dimension of the vector space $H_{N}$ defined in the proof of Theorem 1.5 verifies $\dim(H_{N})\leq\frac{N}{f_{N}g_{N}}$, where $f_{N}$ is given by (4.14), and thus $f_{N}\leq N/(CN^{\beta}+N4^{-\ell_{N}})$. As a consequence of this upper bound on $f_{N}$, for any $b>0$, we may find $a>0$ large enough, and a sequence $(g_{N})$, so that $\dim(H_{N})\leq N/(\ln N)^{b}$. The conclusion follows similarly. Note that it is also possible to pick a larger $\eta_{N}\to 0$ to obtain a vector space $H_{N}$ of smaller dimension. ∎